OpenStudy (anonymous):

Given the general form of the equation of a circle, x2 + y2 + ax + by + c = 0, explain how to (i) write the equation in standard form, and (ii) how to graph the circle in standard form. Also demonstrate (i) and (ii) by providing an example.Can you explain how to graph the equation of a circle. Also demonstrate it by providing an example that have numbers for a, b and c in x2 + y2 + ax + by + c = 0.

OpenStudy (anonymous):

It's easiest to do these in reverse. Let us first start with standard form, expand, and then compare with the original. Here is the equation in standard form for a circle with radius \(r\) located at \((x_0,y_0)\).\[\left(x-x_0\right)^2 + \left(y-y_0\right)^2 = r^2\]This hopefully should be second-nature to you. Let us expand this.\[x^2-2x_0x+x_0^2+y^2-2y_0y+y_0^2=r^2\]Rearranging gives us the following. I have underbraced the parts analogous to the form that you have been given in the problem. \[x^2+y^2\underbrace{-2x_0}_{a}x\underbrace{-2y_0}_{b}y+\underbrace{x_0^2+y_0^2-r^2}_c=0\]Below, I have written exclusively the parts representing a, b, and c. \[a=-2x_0\]\[b=-2y_0\]\[c=x_0^2+y_0^2-r^2\]We have three equations and three unknowns, so we can solve for x_0, y_0, and r. Note: The expression for c was found by substituting the found values of x_0 and y_0.\[x_0 = -\frac{a}{2}\]\[y_0 = -\frac{b}{2}\]\[r^2 = \frac{a^2}{4}+\frac{b^2}{4}-c\]Plugging these values in gives us the answer to part (i).\[\boxed{\left(x+\frac{a}{2}\right)^2+\left(y+\frac{b}{2}\right)^2=\frac{a^2}{4}+\frac{b^2}{4}-c}\]This describes a circle centered at \(\left(-\frac{a}{2},-\frac{b}{2}\right)\) and with radius \(\sqrt{\frac{a^2}{4}+\frac{b^2}{4}-c}\). I trust you can reason out the rest. If you need more help or clarification, just ask.

OpenStudy (anonymous):

wow! Ok- well I got all the way through accept the part explaining how to graph the equation of a circle. Also demonstrate it by providing an example that have numbers for a, b and c in x2 + y2 + ax + by + c = 0.

OpenStudy (anonymous):

Well give me three value for a, b, and c. Then find the center and how big the radius is. That's how you know how to graph it.

OpenStudy (anonymous):

Center is (-a/2, -b/2) Radius is √a2 + b2 -4c/4 if a2 + b2 -4c ≥ 0

OpenStudy (anonymous):

Right, now just give me three numerical values of a, b, and c that work and give me numerical values for the center and radius.

OpenStudy (anonymous):

I have no idea. lol

OpenStudy (anonymous):

thats my problem....the last part of the question.

OpenStudy (anonymous):

Give me three numerical values for a, b, and c of your choosing.

OpenStudy (anonymous):

2, 1,0

OpenStudy (anonymous):

thus where is the center? by the formulas we have?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the center at 0?

OpenStudy (anonymous):

No. Center is (-a/2, -b/2), and you told me a=2 and b=1.

OpenStudy (anonymous):

oh well see, im confused. lol

OpenStudy (anonymous):

if a=2, then what is -a/2?

OpenStudy (anonymous):

a= -2 ?

OpenStudy (anonymous):

no... plug in and evaluate.

OpenStudy (anonymous):

-2/2 = 0 ?! wth

OpenStudy (anonymous):

wait = -1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how about -b/2?

OpenStudy (anonymous):

b= -1

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

b=1, thus, -b/2=?

OpenStudy (anonymous):

-1/2= -0.5

OpenStudy (anonymous):

yep, so our center is (-1,-.5). what's the radius?

OpenStudy (anonymous):

radius is R ??

OpenStudy (anonymous):

yes...what's the numerical value?

OpenStudy (anonymous):

not sure how to get that

OpenStudy (anonymous):

plug in and evaluate the numbers just as before

OpenStudy (anonymous):

i cant do it, im frustrated...just dont know

OpenStudy (anonymous):

\[r=\sqrt{\frac{a^2}{4}+\frac{b^2}{4}-c}\]Every time you see an a, put in your value of a. Every time you see a b, put in your value of b. Every time you see a c, put in your value of c. You'll be left with an expression of all numbers on the right hand side. Simplify it until you get a single number. This is your radius r.

OpenStudy (anonymous):

but what is c?

OpenStudy (anonymous):

you said it was 0

OpenStudy (anonymous):

oh- so i can pick any numbers and it will be right? lol

OpenStudy (anonymous):

as long as they're the same ones you used for the centers...use the same ones you picked before

OpenStudy (anonymous):

That's what it means to given an example.

OpenStudy (anonymous):

*give

OpenStudy (anonymous):

ok i came up with 1.25

OpenStudy (anonymous):

You forgot to square root it. It doesn't come out nice so we can just say \[r=\sqrt{1.25}\]

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

A calculator will give you r=1.11803399.

OpenStudy (anonymous):

ok so how do i put my answer down then for this last question...

OpenStudy (anonymous):

you give your example for a, b, and c, as well as your found center and radius.

OpenStudy (anonymous):

can you show me how to write it though

OpenStudy (anonymous):

HELLO!